Find the points on the ellipse x^2 + 16y^2 = 16 that are furthest
away from the point (0, −1) using the method of Lagrange Multipliers.
1
Expert's answer
2020-04-14T15:05:46-0400

when the distance from the point "(0,-1)" is "d_{(x,y)}" ,


problem is,


"d_{(x,y)}=(x-0)^2+(y-(-1))^2\\
d_{(x,y)}=x^2+(y+1)^2\\
maximize\quad d_{(x,y)}=x^2+(y+1)^2\\
\quad subject\ to,\\
\quad\quad x^2+16y^2=16\\"

Here objective function is "d_{(x,y)} =x^2+(y+1)^2"

and constraint is "C_{(x,y)}=x^2+16y^2-16=0"

so at the maximum point,

"\nabla d_{(x,y)}=\lambda*\nabla C_{(x,y)}\\\quad \lambda\ is\ the\ lagrange\ multiplier"


"\nabla d_{(x,y)}=\begin{bmatrix}2x \\2(y+1) \end{bmatrix}\\"

"\nabla C_{(x,y)}=\begin{bmatrix}2x \\32y\end{bmatrix}\\"

"\begin{bmatrix}2x \\2(y+1) \end{bmatrix}=\lambda\begin{bmatrix}2x \\32y\end{bmatrix}\\"

"2x=\lambda*2x\\
\lambda=1\\
then,\\
 2(y+1)=1*32y\\
y=\frac{1}{15}\\"

"using\ C_{(x,y)},\\
x^2+16*(\frac{1}{15})^2-16=0\\
x=\pm\frac{16}{15}\sqrt{14}\\"

Therefore furthest points are,


"\blue{\bold{(\frac{16}{15}\sqrt{14},\frac{1}{15}) }}\ and\ \blue{ \bold{ (-\frac{16}{15}\sqrt{14},\frac{1}{15})}}\\"



"d_{max}=(\frac{16}{15}\sqrt{14})^2+(\frac{1}{15}+1)^2=\frac{256}{45}"